OpenStudy (anonymous):

Solve 1/2x^2 - 1/2x - 15 = 0

OpenStudy (anonymous):

1/2x^2 - 1/2x - 15 = 0 (1/2x^2 - 1/2x - 15 = 0)2 x^2 - x - 30 = 0 (x-6)(x+5)=0 x=6 and x= -5

OpenStudy (anonymous):

that's the solution kirsten..

OpenStudy (anonymous):

thanks kirsten

OpenStudy (anonymous):

no thank you coolstude haha :)